At 1:00 p.m., a car leaves a city and travels north at a rate of 55 mi/h. An hour later, a second car leaves the city and travels south at a rate of 60 mi/hr. At what time will the two cars be 285 miles apart?
Answer by checkley75(3666)

55X=60(X-1)-285
55X=60X-60-285
55X-60X=-345
-5X=-345
X=-345/-5
X=69 HOURS THEY WILL BE 285 MILES APART WHICH SAYS THEY REACH THIS SEPARATION 2 DAYS & 21 HOURS AFTER THEY LEAVE. THIS PUTS THE CLOCK @ 11 AM.